# Identification of Additive Continuous-time Systems in Open and Closed-loop

### v1

@@ 6.2 Closed-loop Experiment
with sampling period $h=0.05[\textnormal{s}]$. The parameters of interest are $\bm{\beta}^{*}=[0.25,\hskip 3.41418pt0.25,\hskip 3.41418pt3,\hskip 3.41418pt0.01,\hskip 3.41418pt0.025,\hskip 3.41418pt1]^{\top}$. The reference is chosen to be a zero-mean Gaussian white noise signal of unitary variance, and the noise $v(t_{k})$ is a zero-mean white noise Gaussian of variance $0.01$. Fifty different sample sizes are considered, ranging logarithmically from $N=500$ to $N=123036$, each with $300$ Monte Carlo runs with varying reference and noise realizations. Three estimators are tested: the SRIVC estimator (i.e., the direct closed-loop approach, [8]), the CLSRIVC estimator [16], and the proposed estimator. The estimators are all initialized by the same mechanism detailed in Section 6.1, and the relative error bound for the termination rule of each algorithm is set to $10^{-7}$. The SRIVC and CLSRIVC estimators consider a model structure consisting of 4 poles and 2 zeros, and the proposed method exploits the parametrization in (51), which gives $6$ parameters to be computed in total.
Our goal is to compare the parameter vectors of the resulting additive form of each estimator. For this, after computing the estimates using the SRIVC and CLSRIVC methods, we factor the resulting transfer function and obtain their associated additive parameter vector estimate. In Figure 4, we have plotted the mean square error of each parameter for all the sample sizes considered in this study. All three estimators are known to be generically consistent in this setup due to Theorems 10 and 14 of [20], and Theorem 5.5 of this work. However, the proposed method attains lower MSEs compared to the SRIVC and CLSRIVC methods. Similar to Section 6.1, this improvement is due to a more parsimonious model structure that is imposed, as opposed to the unfactored transfer function model descriptions provided by the standard SRIVC and CLSRIVC estimators.
@@ 7 Experimental Validation
Figure 5: Prototype experimental flexible beam. The moving part, i.e., the flexible beam, is equipped with yellow tape. The three actuators are visible on the top part of the image and the five sensors at the bottom. During operation, the beam translates, rotates and exhibits internal flexible behavior.
An open-loop experiment is conducted with a random-phase multisine input of $2$[s] with a flat spectrum between $0.5$[Hz] to $500$[Hz]. Considering the input and output data $\{u(t_{k}),y(t_{k})\}_{k=1}^{8192}$ and the proposed method for the open-loop scenario, the first four modes of the system are estimated (i.e., four second-order systems without zeros), and the result is depicted in Figure 6. The iterations in (32) are initialized with the submodel estimates
@@
These parameters describe a model which deviates on purpose from the expected outcome to illustrate the converging nature of the algorithm. The delay is estimated to be four samples and the converged estimate of the parametric model has parameters
@@ Appendix A Technical Lemma
Lemma A.6 (Generic Non-Singularity).
Consider the $n_{i}$th order, asymptotically stable continuous-time transfer functions $1/A_{i}(p)$, $i=1,\dots,K$, which depend on some parameters $\bm{\nu}_{i}$, and define $A_{i}^{*}(p)$ as the polynomial $A_{i}(p)$ evaluated at $\bm{\nu}_{i}=\bm{\nu}_{i}^{*}$. The polynomials $A_{i}^{*}(p)$, $i=1,\dots,K$, are assumed to be coprime. Furthermore, consider the asymptotically stable discrete-time transfer function $S(q)$, which depends on some parameters $\bm{\xi}$, and define $S^{*}(q)$ as the transfer function $S(q)$ evaluated at $\bm{\xi}=\bm{\xi}^{*}$. Assume that $S(q)$ has at most $m_{s}$ non-minimum phase zeros, and that Assumption 4 holds. If the signal $z(t_{k})$ is stationary and persistently exciting of order no less than $2\sum_{i=1}^{K}n_{i}+m_{s}$, then the following matrix is generically non-singular with respect to $\{\bm{\nu}_{i}\}_{i=1}^{K}$ and $\bm{\xi}$:
$\bm{\Phi}=\mathbb{E}\left\{\begin{bmatrix}S(q)\frac{1}{A_{1}^{2}(p)}\mathbf{z}_{1}(t_{k})\\ S(q)\frac{1}{A_{2}^{2}(p)}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S(q)\frac{1}{A_{K}^{2}(p)}\mathbf{z}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}S^{*}(q)\frac{1}{A_{1}^{*}(p)A_{1}(p)}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{A_{2}^{*}(p)A_{2}(p)}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{A_{K}^{*}(p)A_{K}(p)}\mathbf{z}_{K}(t_{k})\end{bmatrix}^{\top}\right\},$ (52)
@@
    there exist $\{\bm{\nu}_{i}\}_{i=1}^{K}$ and $\bm{\xi}$ vectors that lead to $\bm{\Phi}$ being non-singular.
The first statement follows from applying the same logic as in Lemma 9 of [33] and is therefore omitted. As for the second statement, we will show that setting $S(q)=S^{*}(q)$ and $T_{i}(p)=T_{i}^{*}(p)$ for all $i=1,2,\dots,K$, leads to a non-singular matrix $\bm{\Phi}$. To this end, we define
$\bm{\Phi}^{*}=\mathbb{E}\left\{\begin{bmatrix}S^{*}(q)\frac{1}{[A_{1}^{*}(p)]^{2}}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{[A_{2}^{*}(p)]^{2}}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{[A_{K}^{*}(p)]^{2}}\mathbf{z}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}S^{*}(q)\frac{1}{[A_{1}^{*}(p)]^{2}}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{[A_{2}^{*}(p)]^{2}}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{[A_{K}^{*}(p)]^{2}}\mathbf{z}_{K}(t_{k})\end{bmatrix}^{\top}\right\}.$
@@
$\mathbf{x}^{\top}\bm{\Phi}^{*}\mathbf{x}=\mathbb{E}\left\{\left(\frac{B_{\mathbf{x}}(p)}{\prod_{i=1}^{K}[A_{i}^{*}(p)]^{2}}\tilde{z}(t_{k})\right)^{2}\right\}\geq 0,$ (53)
@@
Now our goal is to show that $B_{\mathbf{x}}(p)\equiv 0$ implies that $Q_{i}(p)\equiv 0$ for all $i=1,\dots,K$, which means that $\mathbf{x}=\mathbf{0}$. First, it is clear that at least two $Q$ polynomials must be non-zero in case the statement were to be disproven. To tackle this case, we argue by contradiction. Suppose that there are at least two polynomials $Q_{k},Q_{l}$ that are non-zero, and without loss of generality [^2] At most one polynomial $Q_{i}$ will have degree $2n_{i}$, due to $n_{i}=m_{i}$ occurring for at most one $i\in\{1,\dots,K\}$. assume that $\text{deg}(Q_{k})<2n_{k}$. Since $B_{\mathbf{x}}(p)\equiv 0$, the following relations between polynomials hold true:
@@
$=-[A_{k}^{*}(p)]^{2}\sum_{\begin{subarray}{c}i=1,\dots,K,\\ i\neq k\end{subarray}}Q_{i}(p)\prod_{\begin{subarray}{c}j=1,\dots,K,\\ j\neq i,k\end{subarray}}[A_{j}^{*}(p)]^{2}$ (61)
@@
Thus, the polynomial in (56) must have the same zeros as $[A_{k}^{*}(p)]^{2}$, counting their multiplicity. These zeros cannot be accounted for in $\prod_{\begin{subarray}{c}j=1,\dots,K,\\ j\neq k\end{subarray}}[A_{j}^{*}(p)]^{2}$ since the $A_{j}^{*}$ polynomials are jointly coprime, and neither in $Q_{k}(p)$, since $\text{deg}(Q_{k})<\text{deg}([A_{k}^{*}(p)]^{2})=2n_{k}$. We conclude that such $Q_{k},Q_{l}$ non-zero polynomials that satisfy the equality in (56) cannot exist. Hence, by contradiction, we have found that all $Q$ polynomials must be zero for $B_{\mathbf{z}}(p)\equiv 0$ to be true, and therefore, $\bm{\Phi}^{*}$ is non-singular.

### v2

@@ 5.2 Closed-loop Experiment
with sampling period $h=0.05[\textnormal{s}]$. The parameters of interest are $\bm{\beta}^{*}=[0.25,\hskip 3.41418pt0.25,\hskip 3.41418pt3,\hskip 3.41418pt0.01,\hskip 3.41418pt0.025,\hskip 3.41418pt1]^{\top}$. The reference is chosen to be a zero-mean Gaussian white noise signal of unitary variance, and the noise $v(t_{k})$ is a zero-mean white noise Gaussian of variance $0.01$. Fifty different sample sizes are considered, ranging logarithmically from $N=500$ to $N=123036$, each with $300$ Monte Carlo runs with varying reference and noise realizations. Three estimators are tested: the SRIVC estimator (i.e., the direct closed-loop approach, [9]), the CLSRIVC estimator [17], and the proposed estimator for closed-loop identification, i.e., (29) with instrument vector described by (20). The estimators are all initialized by the same mechanism detailed in Section 5.1, and the relative error bound for the termination rule of each algorithm is set to $10^{-7}$. The SRIVC and CLSRIVC estimators consider a model structure consisting of 4 poles and 2 zeros, and the proposed method exploits the parametrization in (46), which gives $6$ parameters to be computed in total.
Our goal is to compare the parameter vectors of the resulting additive form of each estimator. For this, after computing the estimates using the SRIVC and CLSRIVC methods, we factor the resulting transfer function and obtain their associated additive parameter vector estimate. In Figure 4, we have plotted the mean square error of each parameter for all the sample sizes considered in this study. All three estimators are known to be generically consistent in this setup due to Theorems 11 and 15 of [22], and Theorem 4.5 of this work. However, the proposed method attains lower MSEs compared to the SRIVC and CLSRIVC methods. Similar to Section 5.1, this improvement is due to a more parsimonious model structure that is imposed: the proposed additive estimator only requires $6$ parameters to exactly describe the system in (46), as opposed to the unfactored transfer function model descriptions provided by the standard SRIVC and CLSRIVC estimators, which require $7$.
@@ 6 Experimental Validation
Figure 5: Prototype experimental flexible beam setup. The moving part is indicated by ⓐ and is suspended by wire flexures ⓑ. The deflection is measured with five contactless fiber optic sensors, of which middle sensor is used ⓒ and the setup is actuated with three current-driven voice coils of which the middle actuator used ⓓ.
An open-loop experiment is conducted with a random-phase multisine input of $2$[s] with a flat spectrum between $0.5$[Hz] to $500$[Hz]. We test the performance of the proposed method for the open-loop scenario, i.e., (29) with instrument vector given by (10). Considering the input and output data $\{u(t_{k}),y(t_{k})\}_{k=1}^{8192}$, the first four modes of the system are estimated (i.e., four second-order systems without zeros), and the result is depicted in Figure 6. The iterations in (29) are initialized with the submodel estimates
@@
These parameters describe a model which deviates on purpose from the expected outcome to illustrate the converging nature of the algorithm and its robustness against poor initialization. The delay is estimated to be four samples and the converged estimate of the parametric model has parameters
@@ Appendix A Technical Lemma
Lemma A.6 (Generic Nonsingularity).
Consider the $n_{i}$-th order, asymptotically stable continuous-time transfer functions $1/A_{i}(p)$, $i=1,\dots,K$, which depend on some parameters $\bm{\nu}_{i}$, and define $A_{i}^{*}(p)$ as the polynomial $A_{i}(p)$ evaluated at $\bm{\nu}_{i}=\bm{\nu}_{i}^{*}$. The polynomials $A_{i}^{*}(p)$, $i=1,\dots,K$, are assumed to be coprime. Furthermore, consider the asymptotically stable discrete-time transfer function $S(q)$, which depends on some parameters $\bm{\xi}$, and define $S^{*}(q)$ as the transfer function $S(q)$ evaluated at $\bm{\xi}=\bm{\xi}^{*}$. Assume that $S(q)$ has at most $m_{s}$ nonminimum phase zeros, and that Assumption 4 holds. If the signal $z(t_{k})$ is quasi-stationary and persistently exciting of order no less than $2\sum_{i=1}^{K}n_{i}+m_{s}$, then the following matrix is generically nonsingular with respect to $\{\bm{\nu}_{i}\}_{i=1}^{K}$ and $\bm{\xi}$:
$\bm{\Phi}=\overline{\mathbb{E}}\left\{\begin{bmatrix}S(q)\frac{1}{A_{1}^{2}(p)}\mathbf{z}_{1}(t_{k})\\ S(q)\frac{1}{A_{2}^{2}(p)}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S(q)\frac{1}{A_{K}^{2}(p)}\mathbf{z}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}S^{*}(q)\frac{1}{A_{1}^{*}(p)A_{1}(p)}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{A_{2}^{*}(p)A_{2}(p)}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{A_{K}^{*}(p)A_{K}(p)}\mathbf{z}_{K}(t_{k})\end{bmatrix}^{\top}\right\},$ (47)
@@
    there exist $\{\bm{\nu}_{i}\}_{i=1}^{K}$ and $\bm{\xi}$ vectors that lead to $\bm{\Phi}$ being nonsingular.
The first statement follows from applying the same logic as in Lemma 9 of [36] and is therefore omitted. As for the second statement, we will show that setting $S(q)=S^{*}(q)$ and $T_{i}(p)=T_{i}^{*}(p)$ for all $i=1,2,\dots,K$, leads to a nonsingular matrix $\bm{\Phi}$. To this end, we define
$\bm{\Phi}^{*}=\overline{\mathbb{E}}\left\{\begin{bmatrix}S^{*}(q)\frac{1}{[A_{1}^{*}(p)]^{2}}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{[A_{2}^{*}(p)]^{2}}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{[A_{K}^{*}(p)]^{2}}\mathbf{z}_{K}(t_{k})\end{bmatrix}\begin{bmatrix}S^{*}(q)\frac{1}{[A_{1}^{*}(p)]^{2}}\mathbf{z}_{1}(t_{k})\\ S^{*}(q)\frac{1}{[A_{2}^{*}(p)]^{2}}\mathbf{z}_{2}(t_{k})\\ \vdots\\ S^{*}(q)\frac{1}{[A_{K}^{*}(p)]^{2}}\mathbf{z}_{K}(t_{k})\end{bmatrix}^{\top}\right\}.$
@@
$\mathbf{x}^{\top}\bm{\Phi}^{*}\mathbf{x}=\overline{\mathbb{E}}\left\{\left(\frac{B_{\mathbf{x}}(p)}{\prod_{i=1}^{K}[A_{i}^{*}(p)]^{2}}\tilde{z}(t_{k})\right)^{2}\right\}\geq 0,$ (48)
@@
Now our goal is to show that $B_{\mathbf{x}}(p)\equiv 0$ implies that $Q_{i}(p)\equiv 0$ for all $i=1,\dots,K$, which means that $\mathbf{x}=\mathbf{0}$. First, it is clear that at least two $Q$ polynomials must be nonzero in case the statement were to be disproven. To tackle this case, we argue by contradiction. Suppose that there are at least two polynomials $Q_{k},Q_{l}$ that are nonzero, and without loss of generality [^3] At most one polynomial $Q_{i}$ will have degree $2n_{i}$, due to $n_{i}=m_{i}$ occurring for at most one $i\in\{1,\dots,K\}$. assume that $\text{deg}(Q_{k})<2n_{k}$. Since $B_{\mathbf{x}}(p)\equiv 0$, the following relations between polynomials hold true:
@@
$=-[A_{k}^{*}(p)]^{2}\sum_{\begin{subarray}{c}i=1,\dots,K,\\ i\neq k\end{subarray}}Q_{i}(p)\prod_{\begin{subarray}{c}j=1,\dots,K,\\ j\neq i,k\end{subarray}}[A_{j}^{*}(p)]^{2}$
@@
Thus, the polynomial in (49) must have the same zeros as $[A_{k}^{*}(p)]^{2}$, counting their multiplicity. These zeros cannot be accounted for in $\prod_{\begin{subarray}{c}j=1,\dots,K,\\ j\neq k\end{subarray}}[A_{j}^{*}(p)]^{2}$ since the $A_{j}^{*}$ polynomials are jointly coprime, and neither in $Q_{k}(p)$, since $\text{deg}(Q_{k})<\text{deg}([A_{k}^{*}(p)]^{2})=2n_{k}$. We conclude that such $Q_{k},Q_{l}$ nonzero polynomials that satisfy the equality in (49) cannot exist. Hence, by contradiction, we have found that all $Q$ polynomials must be zero for $B_{\mathbf{z}}(p)\equiv 0$ to be true, and therefore, $\bm{\Phi}^{*}$ is nonsingular.